# DeFiScope: Detecting Various DeFi Price Manipulations with LLM Reasoning

### v1

@@ 5.2. DeFi Operation Inference
Buy & Sell  (BlockSec, 2024c) I 1) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$ 2) The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases 3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
II 1) The price of $Token_{x}$ in $Pool_{buy}$ increases / The price of $Token_{y}$ in $Pool_{buy}$ decreases/       The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases 2) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$ 3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
Deposit & Borrow  (Immunefi, 2024) III 1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 2) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases 3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
IV 1) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases 2) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
Stake & Claim  (Certik, 2024) V 1) Stake $Token_{x}$ into $Contract_{stake}$ 2) The price of $Token_{y}$ in $Contract_{claim}$ decreases 3) Claim $Token_{y}$ from $Contract_{claim}$
VI 1) The price of $Token_{x}$ in $Contract_{stake}$ increases / The price of $Token_{y}$ in $Contract_{claim}$ decreases 2) Stake $Token_{x}$ into $Contract_{stake}$ 3) Claim $Token_{y}$ from $Contract_{claim}$
Deposit & Withdraw  (SlowMist, 2024a) VII 1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential 2) The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases 3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
VIII 1) The price of $Token_{x}$ in $Contract_{deposit}$ increases / The price of $Token_{y}$ in $Contract_{deposit}$ decreases /       The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases 2) Deposit $Token_{x}$ and get $Token_{y}$ as credential 3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
6. Price Manipulation Detection
@@
The price change information and high-level DeFi operations recovered from §4 and §5, respectively, are finally checked against the detection rules (c.f. Table 1). We analyzed all four specific types of DeFi protocols (c.f. §2.1) and their associated attack instances. Specifically, we examined ElephantMoney (BlockSec, 2024c) for DEX, Cream Finance (Immunefi, 2024) for lending protocols, ATK (Certik, 2024) for staking-based yield farming protocols, and Harvest (SlowMist, 2024a) for deposit-based yield farming protocols, which collectively incurred a loss of \$163.3M.
@@
Based on this in-depth analysis, we identify four attack types targeting different DeFi protocols and their eight generalized attack patterns, as depicted in Table 1. These attack types are Buy & Sell, Deposit & Borrow, Stake & Claim, and Deposit & Withdraw, with each type corresponding to two specific attack patterns. We will discuss these attack types and patterns in more detail.
Buy & Sell. In this type of attack strategy, the attacker primarily profits by first buying $Token_{y}$ with $Token_{x}$ through a swap in $Pool_{buy}$ and then selling $Token_{y}$ for $Token_{z}$ through a swap in $Pool_{sell}$. $Token_{x}$ and $Token_{z}$ can be the same or different tokens. In the attack against ElephantMoney, the attacker first conducted a swap in $Pool_{buy}$ to exchange WBNB for ELEPHANT and then invoked the mint function, which triggered a swap in $Pool_{sell}$ to exchange ELEPHANT for WBNB, resulting in a price increase of ELEPHANT in $Pool_{sell}$. Ultimately, the attacker utilized a reverse swap in $Pool_{sell}$ to obtain WBNB by selling ELEPHANT at the manipulated price. We design Pattern I based on this attack and subsequently generalize it to Pattern II. The major difference between these two patterns is that the token price is manipulated before the first swap in Pattern II, allowing the price of tokens in either $Pool_{buy}$ or $Pool_{sell}$ to be manipulated.
Deposit & Borrow. In this type, the attacker inflates the price of the deposited tokens or deflates the price of the borrowed assets as calculated by $Contract_{borrow}$, bypassing the protective mechanism of over-collateralization, thereby borrowing more assets than the actual value of the collateral. In the Cream Finance incident, the attacker first deposited yUSD as collateral and obtained an equivalent amount of crYUSD as proof of deposit, then inflated the price of yUSD calculated by $Contract_{borrow}$ by transferring a large quantity of yCrv to a specific contract account. Finally, using yUSD as collateral, the attacker borrowed a large amount of assets, which far exceeded the actual value of the deposited yUSD, from $Contract_{borrow}$. We design Pattern III based on this attack and then generalize it to Pattern IV. Pattern IV differs from Pattern III in that the attacker can preemptively increase the price of tokens designated for deposit or decrease the price of assets intended for borrowing as calculated by $Contract_{borrow}$ before the deposit operation. In particular, the motivating example in §2 conforms to Pattern IV.
Stake & Claim. This attack type primarily targets yield-farming protocols that offer staking services. Typically, an attacker first stakes $Token_{x}$ into the application in one transaction. The share ratio of the user is calculated based on the value and quantity of the staked asset in real-time and is stored in the state variables. Then, the attacker decreases the calculated price of $Token_{y}$ in $Contract_{claim}$ and subsequently claims $Token_{y}$ from the contract. $Token_{x}$ and $Token_{y}$ can be the same or different tokens. We derive Pattern V based on the analysis of the attack against ATK. Specifically, in the first transaction, the attacker initially staked ATK into $Contract_{stake}$. Since the staking service required that the ATK be held for 24 hours by the contract account before claiming, the attacker waited for a period and executed the second transaction, exploiting a flash loan to deflate the price of ATK in $Contract_{claim}$, subsequently claiming back an amount of ATK significantly higher than the appropriate quantity. Considering that the attacker can inflate the price of tokens intended for staking beforehand to get an incorrectly calculated share ratio, we further derive Pattern VI from Pattern V.
Deposit & Withdraw. In this attack type, the attacker exploits vulnerabilities in the token pricing mechanism within the deposit or withdrawal contract to conduct price manipulation attacks. We design and generalize Pattern VII based on the Harvest attack. In this hack, the attacker first deposited USDC ($Token_{x}$) in $Contract_{deposit}$ and received fUSDC ($Token_{y}$) as proof. Then, by exchanging USDC for USDT, the price of USDC calculated by $Contract_{withdrawal}$ decreased, and the attacker withdrew an excessive amount of USDC ($Token_{z}$) from $Contract_{withdrawal}$ by burning fUSDC. In this case, $Token_{x}$ and $Token_{z}$ are the same; however, some protocols, such as (Bscscan, 2024), allow different tokens for deposit and withdrawal. Besides deflating the price of tokens to be withdrawn, the attacker can also inflate the price of tokens used for calculating the withdrawal amount, i.e., $Token_{y}$. If the attacker manipulates the token price before depositing, the price of tokens involved in the deposit can also be affected. Based on this assumption, we generalize Pattern VII to create Pattern VIII.
7. Implementation and Evaluation
@@
We implement DeFiScope with about 3,900 lines of Python code. DeFiScope currently supports two blockchains, i.e., Ethereum (Ethereum, 2024) and BSC (Chain, 2024), which account for over 60% of total value locked (TVL) among all the blockchains (DeFiLlama, 2024). To obtain raw data from specific transactions, we utilize blockchain node APIs facilitated by QuickNode (QuickNode, 2024), an external RPC service provider.
Our evaluation aims to answer three research questions (RQs):
@@
Datasets. To address these three RQs, we collect three datasets from real DeFi transactions on Ethereum and BSC to evaluate DeFiScope, as shown in Table 2. Specifically, we use the first dataset, D1, which comprises 95 transactions of real-world price manipulation attacks from 90 DeFi applications, to evaluate both RQ1 and RQ2. To demonstrate DeFiScope’s practicality for RQ3, we use the second dataset, D2, consisting of 968 suspicious transactions collected by our industry partner. Furthermore, to measure DeFiScope’s time overhead when deployed in a realistic setting with both suspicious and benign transactions, we mix the 968 suspicious transactions from D2 with 96,800 benign transactions in the third dataset, D3.
For D1, we scraped data from multiple sources. Initially, we included all 54 price manipulation transactions identified by DeFort (Xie et al., 2024). We also collected all 55 price manipulation attacks documented by the renowned DeFi security GitHub repository, DeFiHackLabs (SunWeb3Sec, 2024), from October 26, 2020, to October 11, 2023. To further expand our dataset, we acquired 31 transactions confirmed as price manipulation attacks from our industry partner. After removing duplicates among the three sources of data, we ultimately obtained 95 price manipulation attacks for D1, which collectively caused \$381.16M in losses on Ethereum and BSC.
We also measure the distribution and monetary losses of the 95 attacks in D1 across our eight attack patterns described in §6. We observe that Pattern I has the highest number of cases, while Pattern III results in the greatest losses. This is largely due to one of the most notorious incidents, CreamFinance (Immunefi, 2024), which falls under this pattern and led to a loss of \$130M. For Patterns V and VIII, there is only one case each: ATK (Certik, 2024) and LUSD (Bscscan, 2024), respectively.
Table 2. The benchmark datasets used for evaluation.
@@
Table 3. Summary of different attack patterns in D1.
@@
D2 comprises 968 suspicious transactions provided by our industry partner, a Web3 security company, which monitors blockchain transactions in real-time and automatically flags transactions that yield significant profits for initiators. All these transactions are sufficiently complex and potentially involve price manipulation. While manually confirming these transactions is labor-intensive and prone to errors, 155 of them have been confirmed to belong to other vulnerabilities, such as Reentrancy (OpenLeverage (Franklin, 2024)), Unverified User Input (YIELD (BlockSec, 2024e)), and Access Control Bugs (SafeMoon (zokyo, 2024)), by developers and industry partners. In this way, we can stress-test DeFiScope’s detection capabilities when mixed with other types of attack transactions.
Furthermore, to assess DeFiScope’s false alarms on benign transactions and to measure its time overhead in a realistic setting, we need to collect a large number of benign transactions, as the most majority of real-world transactions are still benign. In the absence of empirical data on the ratio of suspicious to benign transactions, we adopt a conservative yet reasonable ratio of 1:100, assuming one suspicious transaction for every 100 benign transactions. Thus, with 968 suspicious transactions in D2, we require 96,800 benign transactions for D3. To this end, we randomly sampled these 96,800 transactions from DeFort (Xie et al., 2024)’s dataset of 428,523 benign transactions, which includes 384,143 benign transactions on Ethereum and BSC. At this sample size, we achieve a 99.999% confidence level with a margin of error of 0.625%.
Experimental Setup. All experiments were conducted on a desktop computer running Ubuntu 20.04, powered by an Intel® Xeon® W-2235 CPU (3.80 GHz, 6 cores, and 12 threads) and equipped with 16 GB of memory. For LLMs, we use OpenAI’s GPT family models due to limited data required for fine-tuning, as explained in §4.1. DeFiScope by default uses GPT-3.5-Turbo (GPT-3.5-turbo-1106) for its well-recognized cost-performance balance, but we also conduct an ablation study in §7.2 to test the more advanced GPT-4o (GPT-4o-2024-08-06) for fine-tuning. For the LLM configuration, DeFiScope employs nucleus sampling (Holtzman et al., 2019) with a top-p value of 1 and sets the temperature to 0, ensuring highly deterministic responses for each prompt, although each was run only once.
Table 4. Detection results for 95 ground-truth DeFi price manipulation attacks. DS: DeFiScope, DT: DeFiTainter, DF: DeFort, and DR: DeFiRanger. DeFiScope detected 76 attacks, DeFiTainter detected 34 attacks, DeFort detected 50 attacks, and DeFiRanger detected 49 attacks. Note that DeFiScope failed in 8 cases solely due to ① missing source code and ② compilation errors, which were not caused by DeFiScope’s methodology.
@@ 7.1. RQ1: Detection Effectiveness
To answer RQ1, we evaluate DeFiScope and compare it with three SOTA tools—DeFiTainter (Kong et al., 2023), DeFort (Xie et al., 2024), and DeFiRanger (Wu et al., 2023), using the dataset D1. Because DeFiRanger is not open-source, we use the results reported in their paper for the attacks they evaluated; for other attacks, we re-implemented their approach based on the description (Wu et al., 2023) and conducted our evaluation. Although DeFort is also not open-source, we obtained a copy of its source code from the authors to conduct our experiment.
@@
Table 4 presents the detection results. The first four columns list the name of the victim protocol, the destination chain, the hack date, and the resulting loss, respectively. Note that one protocol may face multiple attacks, so we add numerical suffixes to protocol names to differentiate them. We use ✓ to indicate an attack can be successfully detected by a tool, and ✗ to indicate a detection failure.
@@
DeFiScope can detect most of the price manipulation attacks. It achieves a recall rate of 80%, outperforming all other tools. Overall, DeFiScope detected 76 attacks, followed by DeFort with 50 and DeFiRanger with 49 attacks, respectively, while DeFiTainter detected only 34 attacks. Figure 8 details the performance of each tool on the evaluated DeFi protocols across four application categories using different price models. Compared to other tools, DeFiScope performs the best in every application category. Particularly, it achieves the highest recall, 90.7%, in Token-related protocols. However, DeFiScope yields a low recall rate in Lending-related protocols, though still higher than all existing tools. Through further analysis, we identifies that 3 out of 12 attacks (InverseFinance_2, SanshuInu, and VesperFinance) are cross-transaction attacks that DeFiScope is unable to detect. Additionally, one attack (TIFIToken) involves exploiting a closed-source custom price model, and the InverseFinance_1 attack, a false negative, will be detailed below.
@@
Yet, DeFiScope missed detecting 11 price manipulation attacks. After analyzing each attack, we discovered that 8 out of 11 are cross-transaction attacks (Certik, 2024; Ancilia, 2024b; Phalcon, 2024; Inspex, 2024; Knownsec, 2024; Ancilia, 2024c; BlockSec, 2024d; Finance, 2024), where detection was unsuccessful because DeFiScope is based on analyzing individual transactions. For Zoompro, detection failure occurred because the token involved in the transaction did not adhere to the ERC20 token standard, resulting in the transfer event not being identified. The remaining two attacks, i.e., “IndexedFinance” and “InverseFinance_1” cases, were subject to in-depth analysis. The detection incapability for IndexedFinance was due to its use of an extremely complicated pricing mechanism that involves exponential calculations in its price-related function joinswapExternAmountIn. In the case of InverseFinance, the attacker exploited the flawed price dependency when calculating the price of tokens deposited as collateral, where the collateral price is based on the balance of multiple tokens in the liquidity pool and off-chain price oracles. As both cases require precise quantitative calculations, DeFiScope is limited by the current LLMs’ constrained capacity for scientific computation. A potential solution could be to integrate with Program-Aided Language models (PAL) (Gao et al., 2023), guiding the LLM to generate scripts for necessary calculations and executing them to obtain the result.
@@
Among the remaining 8 undetectable transactions, 3 were not analyzed successfully due to the unavailability of the code of price calculation functions since our method relies on code-level analysis; five cases involved compilation errors during the extraction of price calculation functions using Slither (crytic, 2024), an off-the-shelf static analyzer for Solidity and Vyper, thereby prematurely terminating the detection process. These limitations are not inherent to DeFiScope’s methodology and can be mitigated by automated or semi-automated techniques, e.g., code decompilation (Grech et al., 2019, 2022; Suiche, 2017) and manual intervention.
Figure 8. The categorized detection results on DeFi protocols.
7.2. RQ2: Ablation Study

### v2

@@ V-B DeFi Operation Inference
Buy
Sell
 [31]
I
1) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$
2) The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases
3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
II
1) The price of $Token_{x}$ in $Pool_{buy}$ increases / The price of $Token_{y}$ in $Pool_{buy}$ decreases/
   The price of $Token_{y}$ in $Pool_{sell}$ increases / The price of $Token_{z}$ in $Pool_{sell}$ decreases
2) Swap $Token_{x}$ to $Token_{y}$ through $Pool_{buy}$
3) Swap $Token_{y}$ to $Token_{z}$ through $Pool_{sell}$
Deposit
Borrow
 [32]
III
1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential
2) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases
3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
IV
1) The price of $Token_{x}$ in $Contract_{borrow}$ increases / The price of $Token_{z}$ in $Contract_{borrow}$ decreases
2) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential
3) Borrow $Token_{z}$ using $Token_{x}$ as collateral from $Contract_{borrow}$
Stake
Claim
 [33]
V
1) Stake $Token_{x}$ into $Contract_{stake}$
2) The price of $Token_{y}$ in $Contract_{claim}$ decreases
3) Claim $Token_{y}$ from $Contract_{claim}$
VI
1) The price of $Token_{x}$ in $Contract_{stake}$ increases / The price of $Token_{y}$ in $Contract_{claim}$ decreases
2) Stake $Token_{x}$ into $Contract_{stake}$
3) Claim $Token_{y}$ from $Contract_{claim}$
Deposit
Withdraw
 [34]
VII
1) Deposit $Token_{x}$ into $Contract_{deposit}$ and get $Token_{y}$ as credential
2) The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases
3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
VIII
1) The price of $Token_{x}$ in $Contract_{deposit}$ increases / The price of $Token_{y}$ in $Contract_{deposit}$ decreases /
   The price of $Token_{y}$ in $Contract_{withdrawal}$ increases / The price of $Token_{z}$ in $Contract_{withdrawal}$ decreases
2) Deposit $Token_{x}$ and get $Token_{y}$ as credential
3) Withdraw $Token_{z}$ from $Contract_{withdrawal}$ by burning $Token_{y}$
Based on our in-depth analysis of the top-30 DeFi apps, we identify six DeFi operations that need to be recovered: 1) Swap, primarily from DEXs such as Uniswap; 2) Deposit, defined by yield-farming and lending apps such as AAVE and Pendle; 3) Withdraw, mainly from yield-farming apps; 4) Borrow, based on lending protocols like Compound; 5) Stake and 6) Claim, both predominantly from yield-farming apps like Convex Finance.
Swap involves a user exchanging one token ($Token_{in}$) for another ($Token_{out}$) from a liquidity pool. Relevant contract accounts, excluding user-controlled ones, are noted as liquidity pools for price trend analysis when source code is unavailable. In a Deposit, a user transfers a token ($Token_{deposit}$) to a yield-farming or lending protocol and receives a proof token ($Token_{proof}$) through minting. Withdraw occurs when a user retrieves tokens ($Token_{withdraw}$) from a protocol by burning a proof token ($Token_{proof}$). Borrow refers to a DeFi operation where a borrower receives a token ($Token_{borrow}$) and incurs a debt evidenced by another token ($Token_{debt}$), issued through minting. Stake happens in yield-farming protocols like Convex Finance [35], where users deposit tokens and can earn rewards without receiving a minted proof token. Claim enables users to retrieve staked tokens ($Token_{claim}$) and bonuses without burning any tokens, unlike the Withdraw operation which requires burning tokens to retrieve assets.
An Operation Recovering Example. To recover DeFi operations from the TG, we design and employ a search algorithm based on directed graphs for each operation. Fig. 4 illustrates that, to recover a Swap operation, we use the depth-first search algorithm to identify cycles that both start and end at user-controlled accounts. These cycles must satisfy three constraints: (i) the transfer action of each edge must be transferring token; (ii) the time index of each edge must be monotonically increasing; and (iii) the token involved in the first transfer action must differ from the token in the last transfer action. Therefore, from the well-constructed TG, we can recover the Swap operation expressed as $UC\rightarrow T_{1}\rightarrow CA_{1}\rightarrow T_{4}\rightarrow CA_{2}\rightarrow T_{5}\rightarrow CA_{3}\rightarrow T_{7}\rightarrow UC$, and we label $CA_{1}$, $CA_{2}$, and $CA_{3}$ as liquidity pools.
VI Price Manipulation Detection
@@
The price change information and high-level DeFi operations recovered from §IV and §V, respectively, are finally checked against the detection rules (c.f. Table I). We analyze four prevalent types of DeFi protocols and their associated attack instances. Specifically, we examine ElephantMoney [31] for DEX, Cream Finance [32] for lending protocols, ATK [33] for staking-based yield farming protocols, and Harvest [34] for deposit-based yield farming protocols, which collectively incurred a loss of \$163.3M.
@@
Based on this in-depth analysis, we identify four attack types targeting different DeFi protocols and their eight generalized attack patterns, as depicted in Table I. These attack types are Buy & Sell, Deposit & Borrow, Stake & Claim, and Deposit & Withdraw, with each type corresponding to two specific attack patterns. More details about these attack types and patterns are available in our supplementary material [28].
@@
We implement DeFiScope with about 3,900 lines of Python code. DeFiScope currently supports two blockchains, i.e., Ethereum [36] and BSC [37], which account for over 60% of total value locked (TVL) among all the blockchains [38]. To obtain raw data from specific transactions, we utilize blockchain node APIs facilitated by QuickNode [39], an external RPC service provider.
VII Evaluation
We aims to evaluate the following research questions (RQs):
@@ VII Evaluation
Datasets. To address these RQs, we collect three datasets from real DeFi transactions on Ethereum and BSC to evaluate DeFiScope, as shown in Table II. Specifically, we use the first dataset D1, which comprises 95 transactions of real-world price manipulation attacks from 90 DeFi applications, to evaluate both RQ1 and RQ2. To demonstrate DeFiScope’s practicality for RQ3, we use the second dataset D2, consisting of 968 suspicious transactions collected by our industry partner. Furthermore, to measure DeFiScope’s time overhead when deployed in a realistic setting with both suspicious and benign transactions, we mix 968 suspicious transactions from D2 with 96,800 benign transactions in the third dataset D3.
For D1, we scraped data from multiple sources. Initially, we included all 54 price manipulation transactions identified by DeFort [5]. We also collected all 55 price manipulation attacks documented by the renowned DeFi security GitHub repository, DeFiHackLabs [40], from October 26, 2020, to October 11, 2023. To further expand our dataset, we acquired 31 transactions confirmed as price manipulation attacks from our industry partner. After removing duplicates among the three sources of data, we ultimately obtained 95 price manipulation attacks for D1, which collectively caused \$381.16M in losses on Ethereum and BSC. We also measure the distribution and monetary losses of the 95 attacks in D1 across our eight attack patterns described in §VI, as shown in Table III.
D2 comprises 968 suspicious transactions provided by our industry partner, a Web3 security company, which monitors blockchain transactions in real-time and automatically flags transactions that yield significant profits for initiators. All these transactions are sufficiently complex and potentially involve price manipulation. While manually confirming these transactions is labor-intensive and prone to errors, 155 of them have been confirmed to belong to other vulnerabilities, such as Reentrancy (OpenLeverage [41]), Unverified User Input (YIELD [42]), and Access Control Bugs (SafeMoon [43]), by developers and industry partners. In this way, we can stress-test DeFiScope’s detection capabilities when mixed with other types of attack transactions.
Furthermore, to assess DeFiScope’s false alarms on benign transactions and to measure its time overhead in a realistic setting, we need to collect a large number of benign transactions, as the most majority of real-world transactions are still benign. In the absence of empirical data on the ratio of suspicious to benign transactions, we adopt a conservative yet reasonable ratio of 1:100, assuming one suspicious transaction for every 100 benign transactions. Thus, with 968 suspicious transactions in D2, we require 96,800 benign transactions for D3. To this end, we randomly sampled these 96,800 transactions from DeFort [5]’s dataset of 428,523 benign transactions, which includes 384,143 benign transactions on Ethereum and BSC. At this sample size, we achieve a 99.999% confidence level with a margin of error of 0.625%.
TABLE II: The benchmark datasets used for evaluation.
@@
TABLE III: Summary of different attack patterns in D1.
@@
Minimized Data Leakage. DeFiScope utilizes LLMs to assess token price changes solely by using price calculation functions and token balance changes, without any knowledge of attack transactions. Even if related attacks appeared in pre-training, evaluating price change tendencies in this manner falls outside its training data, resulting in no or minimized data leakage.
Experimental Setup. All experiments were conducted on a desktop computer running Ubuntu 20.04, powered by an Intel® Xeon® W-2235 CPU (3.80 GHz, 6 cores, and 12 threads) and equipped with 16 GB of memory. For LLMs, we by default use OpenAI’s GPT family models but also explore open-source SFT model tuning in §VIII, as explained in §IV-A. DeFiScope by default uses GPT-3.5-Turbo (GPT-3.5-turbo-1106) for its well-recognized cost-performance balance, but we also conduct an ablation study in §VII-B to test the more advanced GPT-4o (GPT-4o-2024-08-06) for fine-tuning. For the LLM configuration, DeFiScope employs nucleus sampling [44] with a top-p value of 1 and sets the temperature to 0, ensuring highly deterministic responses for each prompt, although each was run only once.
TABLE IV: Detection results for 95 ground-truth DeFi price manipulation attacks.
@@ VII-A RQ1: Detection Effectiveness
To answer RQ1, we evaluate DeFiScope and compare it with three SOTA tools, DeFiTainter [7], DeFort [5], and DeFiRanger [4], specialized in detecting price manipulations using the dataset D1. Because DeFiRanger is not open-source, we use the results reported in their paper for the attacks they evaluated; for other attacks, we re-implemented their approach based on the description [4] and conducted our evaluation. Although DeFort is also not open-source, we obtained a copy of code from its authors to run experiments. FlashSyn [45] and SMARTCAT [46] were excluded as baselines because FlashSyn targets broader attacks beyond price manipulations and lacks guidelines for detecting new attacks, while SMARTCAT was unpublished and not open-source at the time of evaluation.
@@
Table IV presents the detection results. The first four columns list the name of the victim protocol, chain deployed, hack date, and resulting loss, respectively. DS, DT, DF and DR denote DeFiScope, DeFiTainter, DeFort and DeFiRanger, respectively. Note that one protocol may face multiple attacks, so we add numerical suffixes to protocol names to differentiate them. We use ✓ to indicate an attack can be successfully detected by a tool, and ✗ to indicate a detection failure.
@@
DeFiScope can detect most of the price manipulation attacks. It achieves a recall rate of 80%, outperforming all other tools. Overall, DeFiScope detected 76 attacks, followed by DeFort with 50 and DeFiRanger with 49 attacks, respectively, while DeFiTainter detected only 34 attacks. In total, DeFiScope failed to detect 19 cases. We manually investigated their root cause and found: (a) 11 cases should be detectable but were missed by DeFiScope due to the limitation of LLM reasoning, non-ERC20 token and the restriction to only single-transaction analysis, and (b) 8 cases cannot be detected because they do not satisfy the source code requirement of DeFiScope due to either ① missing source code or ② compilation errors, as marked in Table IV. Moreover, the statistical significance of DeFiScope’s performance improvements over each baseline is confirmed by the McNemar Test [47], with all the resulting $p$-values being below 0.05 (see details in  [28]). Fig. 5 details the performance of each tool on the evaluated DeFi protocols across four application categories using different price models. Compared to other tools, DeFiScope performs the best in every application category. Particularly, it achieves the highest recall, 90.7%, in the Token category. However, DeFiScope yields a low recall rate in Lending-related protocols, though still higher than all existing tools. Through further analysis, we identifies that 3 out of 12 attacks (InverseFinance_2, SanshuInu, and VesperFinance) are cross-transaction attacks that DeFiScope is unable to detect. Additionally, one attack (TIFIToken) involves exploiting a closed-source custom price model, and the InverseFinance_1 attack, a false negative, will be detailed below.
@@
After analyzing 11 detectable but missed attacks, we discovered that 8 out of 11 are cross-transaction attacks [33, 48, 49, 50, 51, 52, 53, 54], where detection was unsuccessful because DeFiScope is based on analyzing individual transactions. For Zoompro, detection failure occurred because the token involved in the transaction did not adhere to the ERC20 token standard, resulting in the transfer event not being identified. The remaining two attacks, i.e., “IndexedFinance” and “InverseFinance_1” cases, were subject to in-depth analysis. The detection incapability for IndexedFinance was due to its use of an extremely complicated pricing mechanism that involves exponential calculations in its price-related function joinswapExternAmountIn. In the case of InverseFinance, the attacker exploited the flawed price dependency when calculating the price of tokens deposited as collateral, where the collateral price is based on the balance of multiple tokens in the liquidity pool and off-chain price oracles. As both cases require precise quantitative calculations, DeFiScope is limited by the current LLMs’ constrained capacity for scientific computation. A potential solution could be to integrate with Program-Aided Language models (PAL) [55], guiding the LLM to generate scripts for necessary calculations and executing them to obtain the result.
@@
Among the 8 undetectable transactions, 3 were not analyzed successfully due to the unavailability of the code of price calculation functions since our method relies on code-level analysis; five cases involved compilation errors during the extraction of price calculation functions using Slither [56], an off-the-shelf static analyzer for Solidity and Vyper, thereby prematurely terminating the detection process. These limitations are not inherent to DeFiScope’s methodology and can be mitigated by automated or semi-automated techniques, e.g., code decompilation [57, 58, 59] and manual intervention.
Figure 5: The categorized detection results on DeFi protocols.
VII-B RQ2: Ablation Study